OpenStudy (anonymous):

What is the solution to the rational equation 4/3 + 2/x+6 = 18/3x+18

OpenStudy (anonymous):

4/3 + 2/x+6 = 18/3x+18 2/x-6/x=18-4/3 -4/x=50/3 x=(-4)*3/50 x=-6/25

OpenStudy (anonymous):

\[\frac{4}{3}+\frac{2}{x+6}=\frac{18}{3x+18}\]like that?

OpenStudy (anonymous):

Possible Answers: x = –11 x = –3 x = 3 x = 11

OpenStudy (anonymous):

my god

OpenStudy (anonymous):

Its multiple choice

OpenStudy (anonymous):

i am more interested in what the actual question was then the possible answers, but if you have a list of answers then just check to see which one is correct by substituting the number for x and seeing if it works if you want the method, let me know if the question i wrote is correct

OpenStudy (anonymous):

\[4/3=4/(x+6) x=-3 \]

OpenStudy (anonymous):

sorry x=-3

OpenStudy (anonymous):

Thank You